In x-y plane the position coordinates of a particle as a function of time t a is given as x=3t-4t^2 and y=4t+2t^2. The inetial velocity of particle makes an angle with x axis ?

Grade:12th pass

1 Answers

Manas Shukla
7 years ago
We know that
 \frac{\mathrm{d} }{\mathrm{d} x} (position function) = velocity function
So
v_{x} = 3-8t
and
v_{y} = 4+4t
So we know initial velocity
velocity in x direction = 3
velocity in y direction = 4
We know 
\tan \theta = \frac{v_{y}}{v_{x}}
So
\theta = \tan^{-1}\frac{4}{3}